# Youngjoon Lee1, Jinu Gong2, Joonhyuk Kang1 1School of Electrical Engineering, KAIST, South Korea

### v2

@@ IV-B2 Impact of Different Compromise Fractions
In this experiment, we evaluate the plugin’s effectiveness across different fractions of compromised devices. For TMP attacks, our plugin shows remarkable effectiveness across all FL methods as shown in Fig. LABEL:fig:plot2. Initially, vanilla FedAvg’s accuracy drops sharply from 80.06% to 19.47% without plugin as $p$ increases to 0.3. Yet, with plugin applied, FedAvg achieves accuracy above 89.68% even at $p=0.3$. Furthermore, similar improvements are observed in FedProx, FedDyn, and FedRS, where the plugin-enhanced versions maintain accuracy between 85-90% across all compromise ratios. Consequently, these results indicate that our plugin effectively neutralizes targeted attacks by identifying and filtering malicious updates.
Meanwhile, under UMP attacks, the plugin demonstrates strong defense capabilities despite the more challenging attack scenario. Specifically, all methods suffer severe degradation without plugin, dropping to around 17-19% accuracy at $p=0.1$. Yet, FedAvg, FedProx, FedDyn, and FedRS with plugin maintain accuracy between 65-70% even at $p=0.3$. Nevertheless, FedSAM and FedSpeed show limited improvement with plugin under UMP attacks, maintaining only 25-30% accuracy at higher compromise ratios. Indeed, such performance pattern suggests that untargeted attacks pose a greater challenge due to their random nature of perturbations.
Additionally, comparing TMP and UMP scenarios reveals interesting patterns in our plugin’s defensive capabilities. Notably, our plugin provides stronger protection against TMP attacks, with higher accuracy compared to UMP scenarios. Subsequently, FedAvg, FedProx, FedDyn, and FedRS consistently show the best plugin-enhanced performance in both attack types, maintaining above 85% accuracy in TMP and 65% in UMP at $p=0.3$. As $p$ increases, the performance gap between vanilla and plugin-enhanced versions widens, reaching maximum difference at $p=0.3$. Thus, our plugin becomes increasingly crucial as the threat level intensifies.
@@ IV-B3 Impact of Plugin with Different Non-IID Degrees
Lastly, we check the plugin’s effectiveness improves significantly as data heterogeneity decreases ($\alpha$ increases), under TMP attacks at $p=0.3$. Notably, all vanilla methods remain at 20% accuracy regardless of $\alpha$ as shown in Fig. LABEL:fig:plot3. In contrast, plugin-enhanced versions show dramatic improvement beyond $\log\alpha=1$. FedAvg, FedProx, FedDyn, and FedRS with plugin achieve 85-90% accuracy at $\log\alpha=1$. Meanwhile, FedSAM and FedSpeed with plugin reach 70-80% accuracy. Furthermore, this trend indicates that our plugin’s defense mechanism becomes more effective usually as data distributions become more homogeneous.
Moreover, impact of data heterogeneity is more pronounced in untargeted model poisoning scenarios. Specifically, vanilla methods consistently show poor performance around 17-20% across all $\alpha$ values. Additionally, plugin-enhanced methods demonstrate gradual improvement as $\log\alpha$ increases from -1 to 2. FedAvg, FedProx, FedDyn and FedRS with plugin reach 74-75% accuracy at $\log\alpha=2$. Subsequently, FedSAM and FedSpeed follows with 64-65% accuracy. Thus, these results suggest that untargeted attacks remain challenging even with reduced data heterogeneity. Particularly, in highly heterogeneous settings ($\log\alpha<0$), even plugin-enhanced methods struggle to maintain performance above 30%.
Furthermore, performance gap between TMP and UMP scenarios reveals interesting insights about our plugin’s behavior. Consequently, at high $\log\alpha$ values, the gap between TMP and UMP performance is approximately 15-20%. As a result, this consistent difference suggests inherent limitations in defending against untargeted attacks. Nevertheless, steady improvement in both scenarios as $\alpha$ increases demonstrates our plugin’s adaptability. Therefore, while data heterogeneity affects plugin performance, it maintains significant defensive capabilities even under challenging conditions. Ultimately, these findings highlight the importance of considering data distribution characteristics when deploying Byzantine-resilient FL systems.
@@ V Conclusion
In this paper, we propose a intuitive plugin-based approach for achieving Byzantine resilience in heterogeneity-aware FL techniques. Specifically, our solution can be applied seamlessly without modifying core FL algorithms. Through comprehensive experiments on medical image classification task, we validate our proposed approach. Indeed, our virtual data-driven consistency scoring effectively identifies and filters out malicious updates. Furthermore, the plugin maintains high accuracy while providing robust defense. As a result, we achieve significant performance improvements compared to vanilla FL methods. Thus, our plugin offers a practical solution for securing FL systems against Byzantine attacks.

### v3

@@ IV-B2 Impact of Different Compromise Fractions
To check the effect of the proposed plugin, we evaluate its performance under varying fractions of compromised devices. As shown in the top of Fig. LABEL:fig:plot2, our plugin shows remarkable effectiveness across all FL methods under targeted attack. Specifically, vanilla FedAvg accuracy drops from 80.06% to 19.47% as $p$ rises to 0.3, reflecting extreme vulnerability. However, with the plugin enabled, FedAvg maintains above 89% even at the highest compromise level by filtering not only Byzantine updates but also unhelpful local updates. Similar improvements appear in FedProx, FedDyn, and FedRS, which sustain 85–90% accuracy across all values of $p$. The results indicate that the proposed plugin’s ability to detect and filter malicious updates under targeted attacks.
However, as shown in the bottom of Fig. LABEL:fig:plot2, training under untargeted attacks remains more challenging due to their random nature. All vanilla FL methods collapse below 20% accuracy even when $p=0.1$, highlighting their vulnerability. In contrast, plugin-attached FedAvg, FedProx, FedDyn, and FedRS achieve 65–70% accuracy at $p=0.3$, showing considerable gains. As $p$ increases, the gap between vanilla and plugin-attached versions widens markedly, emphasizing the plugin’s growing importance. Thus, our plugin provides meaningful filtering under both attack types and is most effective against targeted attacks.
@@ IV-B3 Impact of Plugin with Different Non-IID Degrees
Finally, we analyze how varying data heterogeneity affects the plugin’s filtering capabilities under Byzantine attacks with $p=0.3$. As shown in the top of Fig. LABEL:fig:plot3, all vanilla FL methods remain stuck near 20% accuracy regardless of $\log\alpha$. However, plugin-attached FedAvg, FedProx, FedDyn, and FedRS achieve 85–90% at $\log\alpha=1$ under targeted attacks, demonstrating substantial gains. FedSAM and FedSpeed follow similar trends, reaching 70–80% under the same conditions. Hence, plugin effectiveness improves as data distributions become more homogeneous and stable. This relationship highlights the interplay between heterogeneity levels and Byzantine resilience degree.
Under untargeted attacks, as shown in the bottom of Fig. LABEL:fig:plot3, the impact of heterogeneity is even more pronounced across all FL methods. In particular, vanilla FL methods remain weak around 17–20% accuracy for all $\alpha$ values, confirming their inherent limitations. Meanwhile, plugin-attached FL methods steadily improve with increasing $\log\alpha$, showing adaptability to distributional shifts. At $\log\alpha=2$, FedAvg, FedProx, FedDyn, and FedRS reach 74–75%, while FedSAM and FedSpeed attain 64–65%. However, when heterogeneity is high ($\log\alpha<0$), even FL methods with plugin struggle to exceed 30%, indicating persistent challenges. The performance gap between TMP and UMP scenarios remains about 15–20% at high $\alpha$, reflecting untargeted attack complexity. However, consistent improvement as $\alpha$ grows confirms that our plugin adapts well to heterogeneous FL settings.
@@ V Conclusion
In this work, we propose a intuitive plugin-based approach to embed Byzantine resilience into heterogeneity-aware FL methods. Moreover, our solution can be attached seamlessly without modifying the core of FL methods. Through extensive experiments on blood cell classification task, we validate the effectiveness of the proposed plugin. In particular, the virtual data-driven consistency scoring accurately detects and filters malicious updates. Thus, our plugin offers a practical way to enable FL methods robust to Byzantine attacks.